### v1

@@
2021
@@
[…]
First principles approaches and concepts for electrochemical systems
Ab initio techniques have revolutionized the way how theory can help practitioners to explore critical mechanisms and devise new strategies in discovering and designing materials. Yet, their application to electrochemical systems is still limited. A well known example how novel concepts can boost our ability to perform such studies is the introduction of temperature control into ab initio simulations. The analogous technique to model electrochemical systems – potential control – is just emerging.
In this review, we critically discuss state-of-the-art approaches to describe electrified interfaces between a solid electrode and a liquid electrolyte in realistic environments. By exchanging energy, electronic charge and ions with their environment, electrochemical interfaces are thermodynamically open systems. In addition, on the time and length scales relevant for chemical reactions, large fluctuations of the electrostatic potential and field occur. We systematically discuss the key challenges in including these features in realistic ab initio simulations, as well as the available techniques and approaches to overcome them, in order to facilitate the development and use of these novel techniques to the wider community. These methodological developments provide researchers with a new level of realism to explore fundamental electrochemical mechanisms and reactions from first principles.
Achieving an atomic scale understanding of electrochemical processes is imperative to achieve breakthrough innovations. Examples are new generations of supercapacitors supercap1 ; supercap2 ; supercap3 ; supercap4 ; supercap5 , metal-air-batteries mab1 ; mab2 ; mab3 ; mab4 ; mab5 , transient electronics transient1 ; transient2 ; transient3 ; transient4 , or new concepts in sustainable metallurgy plasma ; direct ; reduction ; reduction2 . First principles electronic structure techniques are the method of choice. While these techniques are routinely employed in many fields, their application to electrochemical systems is still limited.
A key challenge in describing electrochemical systems and reactions by ab initio techniques is the large range of relevant length and time scales magnussen2019 . The three relevant scales are schematically sketched in Fig. 1. The actual technical system typically of interest is a full cell consisting of two charged electrodes connected to a potentiostat and immersed in an electrolyte (macroscopic scale). A characteristic feature of electrochemical systems is that the applied voltage in the macroscopic system does not drop homogeneously but only in a small region near the electrode-interface. Away from the electrodes the potential is constant, i.e., the system in this region is field free and charge neutral and is, unlike a capacitor, not affected by the electrode distance. From a modeling perspective it can thus easily be captured by suitable boundary conditions.
Figure 1: Schematic representation of an electrochemical system at the three major length- and time-scales - macroscopic, mesoscopic and microscopic - together with the respective electrostatic potential perpendicular to the solid/liquid interface. The distribution of key electrochemical quantities such as surface charge, potential and total energy are shown for the microscopic (top-left) and the meso/macro-scale (bottom-right).
The induced potential is localized in a small region near the electrode-electrolyte interface, resulting in rapid changes in the electrostatic potential close to the electrode (mesoscopic scale). The largest drop and thus the highest electric fields are within the inner and outer Helmholtz layers, which extend over a range of about 1 nm. This length scale corresponds to the atomistic scale and can be represented by modern DFT supercell calculations. While the electrochemically and electrocatalytically most relevant region with the largest electric fields fits within the supercell, many of the control-parameters and boundary conditions, such as the local potential, the surface charge or the total energy, are not constant but show large variations. These are schematically sketched as inset in Fig. 1 by their corresponding density of states, each of which exhibits due to its thermal origin a normal (Gaussian) distribution. It is important to note that this is solely related to the small size of conventional DFT supercells. On the meso- or macroscale, these quantities result from averaging over large areas, essentially yielding constant (delta-function like distributions) as shown in the bottom right inset. Conceptually, these averaged macroscopic quantities can be obtained by performing thermodynamically open supercell calculations (to achieve the desired fluctuations) and average over the resulting supercell configurations.
This concept of stochastic sampling of physical quantities in the supercell is common when performing molecular dynamics (MD) simulations at constant temperature nose1 ; nose2 ; berendsen ; nhchain ; bussi1 ; bussi2 . The correct magnitude of fluctuations in the total energy of the supercell is achieved by applying a thermostat. The thermostat resembles an infinite reservoir with which the supercell exchanges energy to keep the average temperature in the supercell at a predefined value. For finite temperature MD simulations the application of thermostats controls both the average temperature and the thermodynamically consistent distribution of energy fluctuations. The fundamental relation between energy dissipation (to control the temperature) and fluctuations (the stochastic noise term describing the energy fluctuations) is given by the dissipation-fluctuation theorem callen .
While temperature control with thermostats is common and a well-established practice for MD simulations, the application of potentiostats in ab initio MD (AIMD) simulations is still rare magnussen2019 . The reason is on one side the difficulty in setting up computational electrodes. On the other side, most computational setups assume constant voltage or constant charge conditions. While these are the appropriate boundary conditions for the mesoscopic/macroscopic case, as shown in Fig. 1, they fail to capture the voltage and charge fluctuations present on the atomistic level.
In this review, we therefore discuss first in Sec. 2 the various electrostatic boundary conditions and their potential shortcomings when applied within a periodic supercell approach. Sec. 3 reviews available designs of computational electrodes, which are used to induce a potential drop across the electrochemical interface between the working electrode and the electrolyte. Sec. 4 focuses on challenges in applying electric fields in the context of electronic structure simulations, such as band alignment or dielectric breakthrough. Sec. 5 discusses how to set-up ab initio calculations with open thermodynamic boundary conditions, which account for the exchange of electrons with an external reservoir and enable potential control. Finally, we point in the conclusions to some of the challenges in modelling electrified solid/liquid interfaces, which remain open.
Multiple further topics related to ab initio simulations of electrochemical solid/liquid interfaces may be of interest, but are not covered in this review. We therefore encourage interested readers to look into the following reviews about implicit solvents rev_reuter , the electric double-layer rev_gross , the use of phase diagrams in modelling corrosion rondinelly2019 , electrocatalysis koper2012 ; alfonso2018 , electrochemical energy conversion and storage Pastor2024 , energy materials applications butler2019 or water at charged interfaces rev_gonella .

### v2

@@
Abstract
Ab initio techniques have revolutionised the way in which theory can help practitioners to explore critical mechanisms that govern reactions or properties, and to develop new strategies for materials discovery and design. Yet, their application to electrochemical systems is still limited, due to the challenges electronic structure calculations face in achieving a realistic description of electrified solid/liquid interfaces including, e.g., potential and pH control or free energies of barrier configurations. A well-known example of how novel concepts can extend the scope of simulations is the development of thermostats, which introduced temperature control to electronic structure Density Functional Theory (DFT) calculations. The analogous technique for modelling electrochemical systems – potential control, inherent to most electrochemical experiments – is just emerging.
In this review, we critically discuss state-of-the-art approaches to describe electrified interfaces between a solid electrode and a liquid electrolyte in realistic environments. By exchanging energy, electronic charge and ions with their environment, electrochemical interfaces are thermodynamically open systems. In addition, large fluctuations of the electrostatic potential and field occur on the time and length scales relevant to chemical reactions. We systematically discuss the key challenges in incorporating these features into realistic ab initio simulations, as well as the available techniques and approaches to overcome them, in order to facilitate the development and use of these novel techniques by the wider community. These methodological developments provide researchers with a new level of realism to explore fundamental electrochemical mechanisms and reactions from first principles.
keywords:
keyword1, Keyword2, Keyword3, Keyword4
Achieving an atomic scale understanding of electrochemical processes is imperative to achieve breakthrough innovations. Examples are new generations of supercapacitors supercap1 ; supercap2 ; supercap3 ; supercap4 ; supercap5 , metal-air-batteries mab1 ; mab2 ; mab3 ; mab4 ; mab5 , transient electronics transient1 ; transient2 ; transient3 ; transient4 , or new concepts in sustainable metallurgy plasma ; direct ; reduction ; reduction2 . First principles electronic structure techniques such as DFT are usually the method of choice for studying processes at the atomic scale in condensed systems. While these techniques are routinely employed in many fields, their application to electrochemical systems is still limited, due to challenges that DFT calculations face in achieving a realistic description of electrified solid/liquid interfaces including, for example, potential control, pH control or free energies of barrier configurations.
Figure 1: Schematic representation of an electrochemical system at the three major length- and time-scales - macroscopic, mesoscopic and microscopic - together with the corresponding electrostatic potential perpendicular to the solid/liquid interface and laterally averaged. Each of the three electrostatic potentials in the microscopic scale image represents a different configuration obtained from an ab initio molecular dynamics simulation. The fluctuation induced distributions of key electrochemical quantities such as surface charge, potential (with $\sigma_{\Phi}$ being the variance in the potential) and temperature are shown for the microscopic (top-left) and the meso/macro-scale (bottom-right). They show the variation of these quantities over the different snapshots.
A key challenge in describing electrochemical systems and reactions by ab initio techniques is the large range of relevant length and time scales magnussen2019 . The three relevant scales are schematically sketched in Fig. 1. The actual technical system typically of interest is a full cell consisting of two charged electrodes connected to a potentiostat and immersed in an electrolyte (macroscopic scale). A characteristic feature of electrochemical systems is that the applied voltage in the macroscopic system does not drop homogeneously but only in a small region near the electrode-electrolyte interface, usually referred to as the electric double layer. The induced potential drop near the electrode-electrolyte interface is highly localized, with the strongest electric fields confined to the inner and outer Helmholtz layers, typically within a $\sim$ 1 nm region. This atomistic-scale region is well-suited for atomic scale modelling using modern DFT supercell calculations (cf. Fig. 1, left panel).
However, while DFT can capture the sharp potential drop at the interface, the gradual exponential decay of the potential into the bulk solution - influenced by solvated ions - belongs to the mesoscopic regime (cf. Fig. 1, center panel). We emphasize, however, that any ions which specifically adsorb at the interface or are important in the context of reactions can be straightforwardly included within a DFT supercell calculation. Away from the electrodes the potential is constant, i.e., the system in this region is field free and charge neutral and is – unlike a capacitor – not affected by the electrode distance. From a modelling perspective it is thus desirable to capture both the meso-scale decay of the potential outside the DFT supercell and the bulk electrolyte region itself by suitable boundary conditions (cf. also section 3).
Therefore, it is a key concept to describe the region of the strongest electric fields, crucial for electrochemical and electrocatalytic processes, within DFT while embedding the supercell into its surrounding mesoscopic environment via boundary conditions. To achieve this goal, one must ensure that the bounded finite system described within the DFT supercell replicates the exact dynamics of the unbounded, real physical system. However, due to the microscopic size of the supercell the impact of the exchange of e.g. charge and heat with the environment at the boundaries is large. As a consequence, control parameters like the local potential $\Phi_{loc}$, the surface charge, and the total energy are not constant as for the macroscopic system, where these quantities are lateral averages over a macroscopic area. In the limited lateral dimensions of a typical DFT supercell, these quantities exhibit significant fluctuations due to the small number of particles involved. This behavior is illustrated in Fig. 1 by the density of states, where the distributions broaden compared to the macroscopic case. Each of these control parameters exhibits a normal (Gaussian) distribution due to its thermal origin.
We emphasize that such thermally induced fluctuations in conventional DFT supercells arise solely from their small size. They do not arise from artifacts at lateral boundaries, as bulk water properties can be ensured through performing careful tests of supercell size convergence. We note, however, that local fluctuations at reactive sites remain relevant regardless of the cell size. While their statistical distribution becomes size-independent for sufficiently large cells, the fluctuations themselves do not disappear. In fact, they can still strongly influence the reactivity and require explicit treatment via ab initio MD (AIMD) or related methods to be resolved accurately. On the meso- or macroscale, these quantities average over large areas, yielding nearly constant values (delta-like distributions, as shown in the inset). Conceptually, these macroscopic averages can be obtained by performing thermodynamically open supercell calculations to capture the fluctuations, followed by averaging over the resulting supercell configurations.
This concept of statistical sampling of physical quantities in the supercell mirrors the use of thermostats in molecular dynamics (MD) simulations nose1 ; nose2 ; berendsen ; nhchain ; bussi1 ; bussi2 , which mimic an infinite heat reservoir by allowing energy exchange with the environment to maintain a constant temperature . In electrochemical simulations, the analogous process involves maintaining a stable electrode potential by controlling charge exchange with a surrounding reservoir. In finite-temperature MD, thermostats regulate both the average temperature and the thermodynamically consistent energy fluctuation distribution. This interplay between energy dissipation (temperature control) and fluctuations is fundamentally governed by the fluctuation-dissipation theorem callen .
While temperature control with thermostats is common and a well-established practice for MD simulations, the application of potentiostats in AIMD simulations is still rare magnussen2019 . The reason is on one side the difficulty in setting up a suitable surrogate model for the electrochemical double layer. On the other side, most computational setups assume constant voltage or constant charge conditions. While these are the appropriate boundary conditions for the mesoscopic/macroscopic case, as shown in Fig. 1, they cannot capture the sizable voltage and charge fluctuations present on the atomistic level. We note, that the local nature of the fluctuations means that they will have a direct impact on reactions (see discussion in Sec. 3), which are also locally confined in space and time amatore2004 ; santos2024 .
In this review, we start with discussing in Sec. 2 available concepts to include the impact of the electrochemical double layer in the DFT supercell. These electrostatic surrogate models are used to induce a potential drop across the electrochemical interface between the working electrode and the electrolyte. Sec. 3 reviews the various electrostatic boundary conditions and their potential shortcomings when applied to the surrogate models discussed in Sec. 2 within a periodic supercell approach. Sec. 4 discusses how to set-up ab initio calculations with open thermodynamic boundary conditions, which account for the exchange of electrons with an external reservoir and enable potential control. Sec. 5 focuses on practical challenges in applying electric fields in the context of electronic structure simulations, such as band alignment or dielectric breakdown. Finally, we point in the conclusions to some of the challenges in modelling electrified solid/liquid interfaces, which remain open.
Further topics are of interest when performing ab initio simulations of electrochemical solid/liquid interfaces, but are not covered in this review. We refer readers to the following reviews about e.g. implicit solvents rev_reuter and electrification schemes in implicit solvation schwarz2020 , the electric double-layer rev_gross , the use of phase diagrams in modelling corrosion rondinelly2019 , electrocatalysis koper2012 ; alfonso2018 , electrochemical energy conversion and storage Pastor2024 , energy materials applications butler2019 or water at charged interfaces rev_gonella .
2 Design of a surrogate model for the electrochemical double layer
To simulate the electrochemical interface in an ab initio supercell, the potential distribution within the simulation cell must replicate that of the real system in this atomistic region, cf. Fig. 2a. In order to achieve this, a charge on the electrode surface or an electric field needs to be applied (Fig. 1). To maintain charge neutrality under periodic boundary conditions, which is crucial to avoid singularities in the electrostatic energy, the surface charge must be compensated. While many DFT codes address this by introducing a spatially homogeneous background charge, this approach is insufficient for electrochemical systems, where realistic charge compensation must reflect the physics of the electrified interface.
Therefore, in order to achieve realistic field distributions, DFT simulations of electrochemical interfaces commonly employ surrogate models for the electrochemical double layer (cf. Fig. 2b). On the one hand, a primary function of the surrogate model is to compensate for the electrode’s surface charge, ensuring charge neutrality of the simulation cell as a whole. On the other hand, beyond merely neutralizing the system, these models also act as boundary conditions, bridging the explicit simulation region with the broader environment of the double layer that lies beyond the DFT supercell. A critical consideration, as outlined in Sec. 1, is that the electric potential difference between the working electrode and the boundary separating the simulation region from the surrogate model cannot remain constant. Due to the thermal motion of ions and molecular dipoles in the double layer, the potential undergoes continuous fluctuations. Consequently, simply matching the average potential at the boundary is insufficient for achieving a realistic description of a highly dynamic and fluctuating system, such as liquid water. Instead, one must also ensure that the distribution of the thermally induced potential fluctuations at the boundary matches that of the (unbounded) physical system.
Figure 2: a) Comparison between the potential in the macroscopic system and the potential obtained in DFT supercell simulations via a surrogate model for the electric double layer. The potentials need to agree within the microscopic electrochemically active region. To ensure charge neutrality, the residual fraction of the double layer charge that is outside of the supercell must be placed on the double layer surrogate model. Therefore, at larger distances both potentials differ by a constant amount. A suitable correction can be performed a posteriori. For the electrostatic potential $\Phi$ we follow the sign convention common to ab initio calculations, but note that this is opposite to the sign convention used in classical electrostatics. b) The active region described within the supercell. The potential distribution of the macroscopic system is replicated via introducing a surrogate model for the electrochemical double layer. The surrogate model carries a net charge for compensating the surface charge on the explicit working electrode (see Fig. 3 for common realizations of the surrogate model). Note that under reactive conditions, the compensating charge on the surrogate model is not constant but needs to be treated dynamically.
This concept can be more easily understood by considering an analogy with phonons. For an accurate simulation of heat transport, the boundary of the simulation region must be transparent to phonons, allowing them to pass freely rather than reflecting them back into the system. If the boundary is not transparent, heat would accumulate or reflect, resulting in artificial thermal behavior.
Similarly, for electrochemical simulations, the boundary must be transparent to potential fluctuations. Any reflection or suppression of these fluctuations would alter the dynamics and energetics within the simulation cell, leading to unrealistic field distributions and possibly erroneous predictions for processes such as charge transfer or adsorption. By ensuring transparency to potential fluctuations at the boundary, the surrogate model not only captures the correct mean potential but also replicates the thermal equilibrium behavior of the electric field. This is essential for accurately modeling the electrochemical interface and the dynamic interactions that define it.
Realizing a suitable surrogate model has been a challenging task. A natural choice would be including an explicit counter electrode as a specific case, in order to mimic the effect of the double layer beyond the supercell. However, most density-functional theory (DFT) simulations for electrode-electrolyte interfaces are performed using supercell calculations in periodic boundary conditions (PBCs) che1 ; filhol2006 ; taylor2006 ; schnur ; gaigeot ; Galli2021 ; cyli . The use of PBCs, which is required in order to describe extended interfaces, strictly enforces (i) simulation cells that are charge neutral in total (due to Gauss’s law in electrostatics) and (ii) electric field distributions that obey the periodicity of the supercell. Moreover, (iii) in standard DFT codes there is only a single Fermi level that is constant throughout the supercell. It is thus explicitly impossible to maintain a working electrode and an explicit counter electrode within the same supercell at different Fermi levels in standard DFT simulations.
Therefore, while it is easy in the case of thermostats or barostats to adjust the control parameter (i.e. the kinetic energy or the cell volume, respectively), this is considerably more difficult in the case of potentiostats (electrode charge). Multiple approaches to apply fields or charges in supercells have thus been explored over the past three decades, to address the restrictions outlined above. Fig. 3 provides a schematic overview of the main computational setups.
One of the earliest approaches was suggested by Neugebauer and Scheffler neugebauer1993 , extending the original dipole correction neugebauer1992adsorbate beyond merely compensating for the difference in surface dipoles when using an asymmetric slab. By deliberately introducing a net dipole in the supercell the dipole correction acts like a periodic capacitor – effectively placing the slab inside a charged capacitor. Since the dipole correction has been implemented in most plane-wave-based DFT codes it provides an easy approach to apply an external potential. Later, Lozovoi and Alavi lozovoi used a symmetric slab setup, cf. Fig. 3a, where the surfaces are charged by introducing a net charge to the system. In order to ensure charge neutrality of the super cell, an equal and opposite homogeneous background charge is used as a computational counter electrode. While this setup can be easily applied in any DFT code, it does not readily provide a direct measure of the electrode potential (as, for example, the dipole correction neugebauer1992adsorbate ): in principle, the electrode potential is given by the derivative $\Phi=\frac{\partial\mathcal{H}}{\partial q}$ of the DFT total energy $\mathcal{H}$ with respect to the charge $q$, which is, however, a notoriously difficult quantity to compute in 3D PBC supercell calculations defect_review ; FNV_PRL . Also, the distribution of the counter-charge as homogeneous background is very different from the distribution of the counter-charges in a realistic electrochemical cell. In the latter, the counter-charges are realized by the double layer, which shows a distribution very different from a spatially homogeneous and constant background.
Figure 3: The various approaches used to achieve a charged (electrified) electrode-electrolyte interface within a DFT supercell. Electrodes are shown as grey rectangles at the edge(s) of the supercell (marked by a dotted line) and the electrolyte as a blue area within it. Ions and charges at the electrodes are shown as spheres containing a plus or minus, while atoms are shown as coloured spheres. The resulting electrostatic potential curve is depicted by the red line. The realization and distribution of the countercharge introduced by each of these methods to keep the total charge of the supercell at zero is shown in green (filled spheres represent explicit ions, hatched green spheres represent virtual charges that are a consequence of the respective method and do not need to be added explicitly, and green shapes represent continuous densities, e.g. introduced by a constant background (jellium) in Fig. a).
Otani et al. therefore proposed to use 2D PBCs to describe the lateral periodicity of the interface, while lifting the PBC parallel to the surface normal via Green’s function techniques otani2006esm ; greens ; otani2013esm2 , cf. Fig. 3b. Alternatively, one may lift the PBCs selectively only for the electric field in the supercell via electric enthalpy-based approaches pasquarello2002 or Berry phase techniques resta and the modern theory of polarization stengel2007 ; stengel2009 , cf. Fig. 3c. Here, the electrostatic potential drop across the supercell is described by a geometric phase, which accumulates when moving from one periodic image of the supercell to the next. Such approaches to describe electric fields in supercells are demanding to implement. Hence, they are available in only some of the $\sim$ 40+ publicly available DFT codes sandia . Moreover, the practical implementations of this approach typically only work for systems with a bandgap. This is due to the challenges in describing fractional occupancies or band crossings using Wannier functions stengel2007 ; marzari1997 . Thus, only systems with semiconductor or insulator electrodes can be conveniently studied.
A continuous variation of the charge on the working electrode is achieved via charged implicit solvents (Fig. 3d) fattebert2002 ; arias2012 ; mathew2014 ; mathew2019 . These methods effortlessly describe the thermodynamic limit, i.e. do not suffer time-sampling limitations, but are currently unable to describe the local potential fluctuations. Furthermore, hybrid solvation schemes have been suggested (Fig. 3e), where a region of explicit electrolyte is complemented by a charged implicit solvent or continuum dielectric filhol2015 ; gross2015 ; goldsmith ; Melander2024 . These implicit solvents are placed at the edge of the explicit electrolyte region. More recently, QM/MM approaches such as DFT-CES (density functional theory with classical explicit solvents) lim2016 ; ringe2022 have extended traditional implicit solvents beyond (modified) Poisson-Boltzmann models. Hybrid approaches of this type require an implementation within the respective electronic structure package.
Minimally invasive approaches working within the constraints of 3D PBC have been therefore exploited, as well. Multiple groups proposed gche1 ; gche2 ; agross2020 ; omranpoor2023 to vary the number of H+ ions contained within the DFT supercell grand-canonically. For the same purpose, Cucinotta et al. cucinotta introduced an intentionally unbalanced number of cations and anions in the electrolyte. These approaches give rise to an uncompensated net-charge. Since the total system is charge neutral, the non-compensated charge accumulates at the interface, cf. Fig. 3f. By construction, this charge can assume only integer values. Even when using rather large supercells, the number of possible charge states at the interface is limited. This limited number of charge states at the interface severely restricts the variability of the ensuing electric fields and, hence, the applicability of techniques to sample thermodynamically open boundary conditions with $\langle\Phi\rangle=\text{const}$.
For this reason, Surendralal et al. surendralal2018 proposed to replace the charged implicit or explicit region by an explicit atomistic counter electrode using pseudo-atoms, i.e. neutral objects with fractional proton numbers shiraishi , to charge the working electrode. The employed setup introduces an explicit counter electrode consisting of Ne pseudo-atoms and a vacuum region into the supercell, and is shown in Fig. 3g. Apart from enabling a continuous variation of the charge within the supercell, this approach has also the advantage, that: (i) it avoids the use of a compensating background, since the overall supercell remains neutral, (ii) all its design elements can be easily realized in a standard DFT code and do not require changes to the code and (iii) it works equally well with metallic and semiconducting electrodes. Moreover, Surendralal et al.’s setup is straightforwardly extended to describe fully open boundary conditions for the electrode charge deissenbeck2023 . Dudzinski et al. dudzinski2023 recently extended the idea of using fractionally charged pseudo-atoms to modify the total charge of the water molecules in an aqueous electrolyte, although at the price of relinquishing control over the distribution of charges within the electrolyte.
Charging the supercell    & Application to solid/liquid interfaces
Method Charging Geometry Boundary Electrode
slab solvent conditions potential
@@ 2 Design of a surrogate model for the electrochemical double layer
[…]
Sheet charge/jellium lozovoi discrete sym vacuum const $D$ yes
[…]
Electrolyte with net charge
[…]
 Environ environ1 ; environ2 ; goldsmith continuous asym hybrid const $D$ yes
 Explicit H+ gche1 ; gche2 ; agross2020 ; omranpoor2023 discrete asym expl const $D$ yes
 Imbalanced ion conc. cucinotta discrete sym expl const $D$ no
 Fract. charged water dudzinski2023 continuous sym expl const $D$ no
[…]
Grand Canonical-DFT arias2017 continuous asym impl const $E$ yes
Hairy Probes cucinotta2024 continuous asym expl const $E$ yes
@@
Table 1: Methods to charge supercells and model electrochemical solid/liquid interfaces, and their key properties/approximations. We emphasize that the terms ”constant electric field” (const E) and ”constant displacement field” (const D) are common names for the underlying concepts resta ; stengel2007 ; stengel2009 ; sprikPRB and should not be taken literally, i.e., they do not imply a uniform local field or a homogeneous surface charge distribution at the electrode interface. Rather, they refer to the macroscopic laterally averaged boundary conditions imposed at the edges of the periodic simulation cell stengel2009 . The local fields and the surface charge distribution at the electrode interface can be highly inhomogeneous and are determined by the charge self-consistency in the DFT calculations when imposing the macroscopic laterally averaged boundary conditions.
Moreover, we note that methods enforcing a constant displacement field $D$ can, in principle, be adapted to impose a constant electric field $E$ (unless $D$ is constrained by system-specific factors, such as an integer number of compensating ions), but this generally requires an additional self-consistency loop.
We emphasize that modelling the electric double layer either as a charged continuum region outside the explicit electrolyte or as an object consisting of actual atoms creates a hydrophobic gap between the surrogate model and the electrolyte, leading to an undesired and non-negligible potential drop within this region deissenbeck2023 . Moreover, this interfacial potential loss at the surrogate model increases the likelihood of reaching a dielectric breakdown yoo2021 as the field strength increases (see also discussion in section 5).
To eliminate the hydrophobic gap at the computational counter electrode and any associated potential loss, in a newer development deissenbeck2023 the setup proposed by Surendralal et al. surendralal2018 has been extended to the use of a fully solvated electrode, cf. Fig. 3h. In this approach a Gaussian counter charge, placed within the aqueous region, is used to charge the working electrode. To maintain the desired bulk water density, a confining force is applied to the water surface facing the periodic image of the working electrode. In the ensuing vacuum region, the dipole correction neugebauer1992adsorbate is employed to remove any electric fields and eliminate spurious interactions between periodic images. On the right side of the cell, the potential is given by the workfunction of the uncharged (backside) surface of the slab facing the vacuum region (as in Fig. 3e and g). This potential can be calculated for a free-standing uncharged slab in vacuum and its absolute value (workfunction) can be straightforwardly computed and used as an absolute potential reference.
While this method requires slight modifications in the DFT code, it enables considerably larger potential drops and realistically strong electric fields at the working electrode/water interface. Moreover, this development allows to apply arbitrarily shaped countercharges. By determining the countercharge distribution self-consistently, such an approach could accurately describe the ionic charge distribution of the electrochemical double layer in that region, in order to realistically capture the diffuse region at the interface.
As a final point, we emphasize that in DFT simulations of electrochemical interfaces, only the near-surface region of the double layer is captured within the supercell, while the gradual potential decay toward the bulk electrolyte remains unresolved (cf. Fig. 2a). This gap can be bridged by applying a posteriori corrections using physically motivated models of the double layer, such as Poisson–Boltzmann distributions or electrostatic models informed by the computed charge distribution near the interface. These approaches are conceptually similar to methods used to correct long-range electrostatic effects in charged supercells defect_review , where the external potential is approximated based on physical constraints.
Table 1 presents a summary of current methods used in DFT calculations to model electrochemical interfaces, highlighting their geometrical set-ups, charging capabilities, as well as their ability to enable electrode potential control. It also indicates the electrostatic boundary conditions commonly used within the respective schemes. These boundary conditions are discussed in detail in the following section.